# Provable Emergence of Deep Neural Collapse and Low-Rank Bias in $L^{2}$-Regularized Nonlinear Networks

### v2

@@ 3.2 Neural collapse yields low-rank bias
$\operatorname{\mathcal{L}}^{l,\mathcal{C}}_{\lambda}(\mathit{\Theta})=\sum_{k=1}^{K}\pi_{k}\operatorname{\ell}\big{(}f_{l}(\mathit{\Theta}_{l},\bar{z}_{k}^{l}),\bar{y}_{k}\big{)}+\lambda\|\mathit{\Theta}\|^{2},\quad\pi_{k}=\frac{N_{k}}{N}.$ (8)
@@
Lemma 3.3 (Centroid-based gradients approximate full gradients for well-clustered data).
@@
$\begin{array}[]{lr}\operatorname{\nabla}_{W_{l}}\operatorname{\mathcal{L}}_{0}(\mathit{\Theta})=\nabla_{W_{l}}\operatorname{\mathcal{L}}_{0}^{j,\mathcal{C}}(\mathit{\Theta})+R(\mathit{\Theta}),\quad\|R(\mathit{\Theta})\|\leq M_{l,j}(\mathit{\Theta},\mathcal{C})\operatorname{\mathrm{WCSS}}_{j}(\mathcal{C})\end{array}$ (9)
@@
Theorem˜3.4 shows that an intermediate neural collapse combined with weight decay regularization leads to approximately low-rank network layers, where the numerical ranks effectively depend on the $\operatorname{\mathrm{TCV}}$ of intermediate layers. Its proof can be found in Section˜A.3. These considerations are directly connected with the notion of deep (or intermediate) neural collapse sukenik2024deep ; papyan2020traces ; rangamani2023feature , whose first characterizing property is recalled below:
Definition 3.5 (Deep Neural Collapse 1).
We say that a neural network $f$ trained on a dataset $\mathcal{X},\mathcal{Y}$ satisfies $\operatorname{\mathrm{DNC1}}$ at layer $l$ if $\operatorname{\mathrm{TCV}}_{K,l}=0$ for $K=c$ and the minimizer in (7) is given by the partition corresponding to the classes.
Notice that the vanishing property of $\operatorname{\mathrm{TCV}}$ in intermediate layers ($\operatorname{\mathrm{DNC1}}$) has already been studied for simplified models in, e.g., Poggio_hidden_nc ; sukenik2024deep ; sukenik2024neural ; jacot2025wide . We therefore notice here the special role of $\operatorname{\mathrm{DNC1}}$ with respect to the other $\mathrm{DNC}$ properties commonly studied, such as the ones presented in sukenik2024deep : the first property alone implies that the stationary points in parameter space must have low-rank, as predicted by Theorem˜3.4.
In the next section, we will prove that, in a fairly general setting, $\operatorname{\mathrm{DNC1}}$ is optimal for all intermediate layers: this means that under a set of natural constraints, a zero total cluster variation for all intermediate layers is globally optimal in the small $\lambda$ regime. This shows, in combination with Theorem˜3.4, that all global minima in this regime are low rank. Moreover, we will show that for almost every initial condition, there is essentially no loss barrier to configurations with $\operatorname{\mathrm{TCV}}=0$ for all $l=1,\dots,L$, shedding new light on the emergence deep neural collapse and, in turn, on the emergence of low-rank bias.
@@ 4 Vanishing $\operatorname{\mathrm{TCV}}$: global optimality of deep-neural collapse
$\mathscr{L}(\mathit{\Theta}):=\begin{cases}+\infty,\quad\text{if}\,\,f(\mathit{\Theta};X)\neq Y,\\ \frac{1}{2}\sum_{l=1}^{L-1}\|W_{l+1}\|^{2},\quad\text{otherwise}.\end{cases}$ (10)
This functional has been studied in a variety of different settings, both linear Ranzato_2021 and nonlinear jacot2023bottleneck ; jacot2023implicit . While in the deep linear case the explicit form of the minima of $\mathscr{L}$ are known to be Shatten quasi-norms Ranzato_2021 ; ongie2022role , in the general case a full characterization is still missing. In (jacot2023implicit, ; jacot2023bottleneck, ), the authors were able to characterize, in the large depth regime, how the minimal value essentially depends on the data $y=f^{*}(x)$ one is trying to interpolate, essentially showing that the first order term scales with $L$ and it is a function bounded between the Jacobian and the bottleneck rank of the function $f^{*}$.
The general intuition is that $\mathscr{L}(\mathit{\Theta})$ represents the loss of interpolating networks in the small $\lambda$ regime, in fact from Scagliotti_2022 , we have that $\lambda^{-1}\mathcal{L}_{\lambda}\underset{\lambda\to 0}{\to}\mathscr{L}(\mathit{\Theta})$ in the $\Gamma$-convergence sense (Definition˜A.1), meaning that minimizers of $\mathcal{L}_{\lambda}$ converge to minimizers of $\mathscr{L}$ as $\lambda$ goes to zero, see Proposition˜A.2. Now, using the observation done in jacot2022featurelearningl2regularizeddnns , we can reparameterize the problem using minimal norm feature matrices in the following way:
$Z^{l+1}=\sigma_{l+1}(W_{l+1}Z^{l})\iff W_{l+1}=\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+},\,\,\text{s.t.}\,\,\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}Z^{l}=\sigma_{l+1}^{-1}(Z^{l+1}),$
where $Z^{l,+}$ denotes the Moore-Penrose pseudoinverse of the feature matrix $Z^{l}$. Notice that the constraint $\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}Z^{l}=\sigma_{l+1}^{-1}(Z^{l+1})$ is effectively a representability condition on the features, equivalent to $W_{l+1}Z^{l}=\sigma_{l+1}^{-1}(Z^{l+1})$. With this new representation, we can study the representation cost as a function of the feature matrices $Z^{l}$, namely:
@@
As proven in jacot2022featurelearningl2regularizeddnns , it is not restrictive to study this reparametrized representation cost in the sense that the local minimizers of (10), subject to the representability constraints (I), are in bijection with those of (11), therefore making them equivalent. The key observation is that the objective in (11) is local, i.e., it is a sum of local objectives $\mathscr{L}_{l}=\mathscr{L}_{l}(Z^{l+1},Z^{l})$ depending just on neighbour representations. Therefore, since $\|W_{1}\|^{2}$ is not present in the regularization, $\mathscr{L}$ depends on $Z^{1}$ just through $\mathscr{L}_{1}(Z^{2},Z^{1})=\frac{1}{2}\|\sigma_{2}^{-1}(Z^{2})Z^{1,+}\|^{2}$, which can be minimized explicitly with respect to $Z^{1}$, as a function of the next representation $Z^{2}$. For this reason, the lack of the first layer in the regularization term is the key to obtaining the results in the following section, as it makes it possible to study recursively the structure of minima layer by layer. To study it, we consider the additional hypothesis that $Y$ is in one-hot encoding form, i.e. $y_{j}=e_{i_{j}}$ for any $j$, where $\{e_{i}\}$ is the canonical basis of $\mathbb{\mathbb{R}}^{c}$. We also point out that, without considering additional constraints, the minima of $\mathscr{L}$ may not exist. To avoid these situations, we study the minima of $\mathscr{L}$ subject to additional constraints:
@@
Consider the set $\mathcal{S}$ consisting of all networks $Z=(Z^{1},\dots,Z^{L})\in\mathbb{R}^{n_{1}\times N}\times\dots\times\mathbb{R}^{n_{L}\times N}$ that satisfy the following conditions (I)-(II)-(III) for all $l=1,\dots,L-1$:
@@
    $\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}Z^{l}=\sigma_{l+1}^{-1}(Z^{l+1})$,
@@
    $s_{i}(Z^{l})\leq Cs_{i}(\sigma_{l+1}^{-1}(Z^{l+1})),\quad\forall i=1,\dots,r_{l}$,
@@
where $C>0$ is arbitrary, $r_{l}$ is the rank of $\sigma_{l+1}^{-1}(Z^{l+1})$, $s_{i}(A)$ denotes the $i$-th singular value of a matrix $A$ and $\sigma_{2},\dots,\sigma_{L}$ is a set of entrywise omeomorphisms with $\sigma_{l}(x)\neq 0$ for all $x\neq 0$.
We highlight here that constraint (II) corresponds to a constraint just on the nonzero singular values of the matrices $Z^{l}$ and $\sigma_{l+1}^{-1}(Z^{l+1})$. Therefore, (II) allows the case in which all feature matrices $Z^{l}$ for $l=1,\dots,L-1$ are full-rank. Moreover, $\mathcal{S}$ is bounded but not compact, making it nontrivial to prove the existence of a minimizer. The following main result on the minimizers of a constrained version of (11) holds:
@@
The global minimizers of the optimization problem
@@
satisfy $\operatorname{\mathrm{DNC1}}$ for all intermediate layers $l=1,\dots,L$. In particular, any minimum point $\bar{Z}$ satisfies
$\bar{Z}^{l}\propto O_{l}\sigma_{l+1}^{-1}(\bar{Z}^{l+1}),\quad O_{l}\in\mathbb{R}^{n_{l}\times n_{l+1}},\quad O_{l}^{\top}O_{l}=I,\quad\forall\,l=1,\dots,L-1$
where the proportionality constant is $C$ in Definition˜4.1.
A rigorous formulation and proof of Theorem˜4.2 can be found in Appendix˜B.
Theorem˜4.2 can be reinterpreted as follows: if we allow the architecture to collapse already at the first layer, that is, if the kernel of $W_{1}$ can be large enough (i.e., $n_{1}\geq N$), then, under (I)(III), it is optimal to collapse all intermediate layers. Notice that the first set of constraints (I) has a crucial role in terms of the representability of the features, and it effectively couples the representations of different layers, as discussed above. The second constraint (II) can be directly related to the boundedness of the nonlinear activations, and it represents exactly the situation for diagonal feature matrices, for which the nonlinearity acts directly on the singular values. The necessity of constraint (II) relies in making it possible to study the alignment of the optimal basis, avoiding any possible singular value escaping to infinity. It is interesting to notice that while in (II) $C$ is an arbitrary positive constant, we observe numerically in Section˜5 that a perfect fit happens for $C=1$, reminiscent of a norm balancing condition on all the layers. The third set (III) is introduced here as it arises naturally in the proof of the theorem as a consequence of (II), and it allows us to rule out the presence of global minima of lower rank. This is because, as we will show in the proof in Appendix˜B, all global minima satisfy the equality constraint. We notice that, as a byproduct of Theorem˜4.2, we know that if we get rid of constraint (III), then all global minimizers either satisfy $\operatorname{\mathrm{DNC1}}$ or some layer embedding has rank strictly smaller than $K$. We expect these minimizers to be exceptional in some sense, as they were proven to be optimal in a very particular setting in sukenik2024neural . Morever, we point out that also the singular values structure is predicted recursively by Theorem˜4.2, which we will numerically show also in Section˜5.
We emphasize that Theorem˜4.2 can be extended to residual networks under a weaker set of assumptions. Indeed, in that case, we can get rid of all constraints, except the natural one in (I). Due to space limitations, we present this result along with its proof in Appendix˜C.
@@
Theorem 4.3 ($\operatorname{\mathrm{DNC1}}$ is reachable for almost every initial condition).
@@
Assume the intermediate widths in $\mathcal{S}$ satisfy the condition $K\leq n_{L}=n_{L-1}=\dots=n_{1}=N$ and that the $\sigma_{l}$’s are analytic. Let $Z(0)\in\mathcal{S}$ be a random initial interpolating condition, i.e., such that $Z^{L}(0)=Y$. Then, almost surely there exists a continuous path $\gamma(t)$ with $\gamma(0)=Z(0)$, $\gamma(1)$ global minimum of $\mathscr{L}$, and such that $\mathscr{L}\circ\gamma$ is continuous Lebesgue-almost everywhere (except for a finite set of times) and for all continuity times $s\geq t$ we have
$\mathscr{L}(\gamma(t))\geq\mathscr{L}(\gamma(s)).$
Moreover, all discontinuities $t_{1},\dots,t_{m}$ are removable and satisfy
$\underset{t\to t_{i}^{-}}{\lim}\mathscr{L}(\gamma(t))=\underset{t\to t_{i}^{+}}{\lim}\mathscr{L}(\gamma(t))>\mathscr{L}(\gamma(t_{i})),\quad\forall i=1,\dots,m.$

### v3

@@ 3.2 Neural Collapse Yields Low-Rank Bias
$\operatorname{\mathcal{L}}^{l,\mathcal{C}}_{\lambda}(\mathit{\Theta})=\sum_{k=1}^{K}\pi_{k}\operatorname{\ell}\big(f_{l}(\mathit{\Theta}_{l},\bar{z}_{k}^{l}),\bar{y}_{k}\big)+\lambda\|\mathit{\Theta}\|^{2},\quad\pi_{k}=\frac{N_{k}}{N}.$
@@
Lemma 3.4 (Centroid-based gradient approximation of full gradient).
@@
$\begin{array}[]{lr}\operatorname{\nabla}_{W_{l}}\operatorname{\mathcal{L}}_{0}(\mathit{\Theta})=\nabla_{W_{l}}\operatorname{\mathcal{L}}_{0}^{j,\mathcal{C}}(\mathit{\Theta})+R(\mathit{\Theta}),\quad\|R(\mathit{\Theta})\|\leq M_{l,j}(\mathit{\Theta},\mathcal{C})\operatorname{\mathrm{WCSS}}_{j}(\mathcal{C})\end{array}$
@@
Proof (sketch).
Here we provide a sketch of the proof and we refer to Appendix A.2 for a detailed argument. Fixed a single data point $(z_{i}^{j},y_{i})$ with $i\in\mathcal{C}_{k}$, we consider the Taylor expansion in $(\bar{z}_{k}^{j},\bar{y}_{k})$ of $\nabla_{W_{l}}\ell(f_{j}(\mathit{\Theta}_{j},z_{i}^{j}),y_{i})$ with Lagrange remainder. Finally, we note that the first-order contributions sum to zero because of the definitions of the centroids, yielding the desired expressions. ∎
@@
Proof (sketch).
Observe that from equation 4, at the stationary point we have
Now, Lemma 3.4 allows us to control the distance of the gradient in the right-hand side of equation 14 from $\mathcal{M}_{K}^{d_{l-1},d_{l}}$. In particular, given $\mathcal{C}\in\cup_{r=1}^{K}\mathcal{P}_{r}$, for any $j=1,\dots,l-1$ we see that
$\min_{Z\in\cup_{r=1}^{K}\mathcal{M}_{r}}\|W^{*}_{l}-Z\|$ $\leq\frac{\|\operatorname{\nabla}_{W_{l}}\operatorname{\mathcal{L}}_{0}(\mathit{\Theta}^{*})-\operatorname{\nabla}_{W_{l}}\operatorname{\mathcal{L}}_{0}^{j,\mathcal{C}}(\mathit{\Theta}^{*})\|}{\lambda}\leq\frac{M_{l,j}(\mathit{\Theta}^{*},\mathcal{C})\operatorname{\mathrm{WCSS}}_{j}(\mathcal{C})}{\lambda},$
where we have used that $\operatorname{\nabla}_{W_{l}}\operatorname{\mathcal{L}}_{0}^{j,\mathcal{C}}(\mathit{\Theta}^{*})/\lambda$ has rank smaller than or equal to the number of families of $\mathcal{C}$. Then, first minimizing over $j=1,\dots,l-1$, second bounding from above $M_{l,j}(\mathit{\Theta}^{*},\mathcal{C})\leq M_{l,j}(\mathit{\Theta}^{*})$ for any $j$, and third minimizing over all the possible partitions $\mathcal{C}\in\cup_{r=1}^{K}\mathcal{P}_{r}$ yield the thesis. ∎
Theorem˜3.5 shows that an intermediate neural collapse, Definition˜3.3, combined with weight decay regularization leads to approximately low-rank network layers, where the numerical ranks effectively depend on the $\operatorname{\mathrm{TCV}}$ of intermediate layers. Its proof can be found in Section˜A.3.
Notice that the vanishing property of $\operatorname{\mathrm{TCV}}$ in intermediate layers ($\operatorname{\mathrm{DNC1}}$) has already been studied for simplified models in, e.g., (Rangamani et al., 2023a; Súkeník et al., 2024b; a; Jacot et al., 2025). We therefore notice here the special role of $\operatorname{\mathrm{DNC1}}$ with respect to the other $\mathrm{DNC}$ properties commonly studied, such as the ones presented in (Súkeník et al., 2024b): the first property alone implies that the stationary points in parameter space must have low-rank, as predicted by Theorem˜3.5.
In the next section, we will prove that, in a fairly general setting, $\operatorname{\mathrm{DNC1}}$ is optimal for all intermediate layers: this means that under a set of natural constraints, a zero total cluster variation for all intermediate layers is globally optimal in the small $\lambda$ regime. This shows, in combination with Theorem˜3.5, that all global minima in this regime are low rank. Moreover, we will show that for almost every initial condition, there is essentially no loss barrier to configurations with $\operatorname{\mathrm{TCV}}=0$ for all $l=1,\dots,L$, shedding new light on the emergence deep neural collapse and, in turn, on the emergence of low-rank bias.
@@ 4 Vanishing $\operatorname{\mathrm{TCV}}$: Global Optimality of Deep-Neural Collapse
$\mathscr{L}(\mathit{\Theta}):=\begin{cases}+\infty,\quad\text{if}\,\,f(\mathit{\Theta};X)\neq Y,\\ \frac{1}{2}\sum_{l=2}^{L-1}\|W_{l+1}\|^{2},\quad\text{otherwise}.\end{cases}$ (9)
This functional has been studied in a variety of different settings, both linear (Dai et al., 2021) and nonlinear (Jacot, 2023a; b). While in the deep linear case the explicit form of the minima of $\mathscr{L}$ are known to be Shatten quasi-norms (Dai et al., 2021; Ongie and Willett, 2022), in the general case a full characterization is still missing. In (Jacot, 2023b; a), the authors were able to characterize, in the large depth regime, how the minimal value essentially depends on the data $y=f^{*}(x)$ one is interpolating, essentially showing that the first order term scales with $L$ and it is a function bounded between the Jacobian and the bottleneck rank of the function $f^{*}$.
The general intuition is that $\mathscr{L}(\mathit{\Theta})$ represents the loss of interpolating networks in the small $\lambda$ regime, in fact from (Scagliotti, 2022), we have that $\lambda^{-1}\mathcal{L}_{\lambda}\underset{\lambda\to 0}{\to}\mathscr{L}(\mathit{\Theta})$ in the $\Gamma$-convergence sense (Definition˜A.1), meaning that minimizers of $\mathcal{L}_{\lambda}$ converge to minimizers of $\mathscr{L}$ as $\lambda$ goes to zero, see Proposition˜A.2. Now, using the observation done in (Jacot et al., 2022), we can reparameterize the weight matrices $\{W_{l}\}$ using the feature matrices in the following way:
$Z^{l+1}=\sigma_{l+1}(W_{l+1}Z^{l})\iff W_{l+1}=\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+},$ (10)
where $Z^{l,+}$ denotes the Moore-Penrose pseudoinverse of the feature matrix $Z^{l}$. If we want to reformulate the representation cost function in equation 9 based on the equivalence discussed in equation 10 we have to additionally take into account for a constraint relating the different intermediate representations, literarily $Z^{l}$ and $Z^{l+1}$ have to satisfy the following equation
$\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}Z^{l}=\sigma_{l+1}^{-1}(Z^{l+1})\qquad\forall l=0,\dots,L-1.$ (11)
Note that the constraint in equation 11 is effectively a representability condition on the features, equivalent to $W_{l+1}Z^{l}=\sigma_{l+1}^{-1}(Z^{l+1})$. Given this new representation of the weight matrices, we can reformulate the representation cost as a function of the feature matrices $Z^{l}$, namely:
@@
As proven in (Jacot et al., 2022), it is not restrictive to study this reparametrized representation cost in the sense that the local minimizers of equation 9, subject to the representability constraints 11, are in bijection with those of equation 12, therefore making them equivalent. The key observation is that the objective in equation 12 is local, i.e., it is a sum of local objectives $\mathscr{L}_{l}=\mathscr{L}_{l}(Z^{l+1},Z^{l})$ depending just on neighbour representations. Therefore, since $\|W_{1}\|^{2}$ is not present in the regularization, $\mathscr{L}$ depends on $Z^{1}$ just through $\mathscr{L}_{1}(Z^{2},Z^{1})=\frac{1}{2}\|\sigma_{2}^{-1}(Z^{2})Z^{1,+}\|^{2}$, which can be minimized explicitly with respect to $Z^{1}$, as a function of the next representation $Z^{2}$. For this reason, the lack of the first layer in the regularization term is the key to obtaining the results in the following section, as it makes it possible to study recursively the structure of minima layer by layer. To study it, we consider the additional hypothesis that $Y$ is in one-hot encoding form, i.e. $y_{j}=e_{i_{j}}$ for any $j$, where $\{e_{i}\}$ is the canonical basis of $\mathbb{\mathbb{R}}^{c}$. We also point out that, without considering additional constraints, the minima of $\mathscr{L}$ may not exist. To avoid these situations, we study the minima of $\mathscr{L}$ subject to additional constraints.
4.1 Characterizing Optimal Representations: Natural Architectural Constraints
Before stating Theorem˜4.2, we introduce three sets of constraints that capture essential properties of realizable representations in actual neural networks. These constraints are a way to encode properties of the underlying architecture in the space of representable intermediate embeddings. In particular, this allows one to view dually the neural network under consideration as a pure constraint on the space of intermediate representations, which forces the embedding of one layer $Z^{l}$ to be linked to that of the embedding before $Z^{l-1}$ through the imposed architecture. We first start with the definition of the constraints.
@@ 4.1 Characterizing Optimal Representations: Natural Architectural Constraints
Consider the set $\mathcal{S}=\mathcal{S}(Z^{1})$ consisting of all networks $Z=(Z^{2},\dots,Z^{L})\in\mathbb{R}^{n_{2}\times N}\times\dots\times\mathbb{R}^{n_{L}\times N}$ that satisfy the following conditions (I)-(II)-(III):
@@
    $\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}Z^{l}=\sigma_{l+1}^{-1}(Z^{l+1}),\quad\forall l=1,\dots,L-1$;
@@
    $s_{i}(Z^{l})\leq C_{l+1}s_{i}(\sigma_{l+1}^{-1}(Z^{l+1})),\quad\forall i=1,\dots,r_{l},\quad\forall l=2,\dots,L-1$,
@@
where $C_{l}>0$ are arbitrary, $r_{l}$ is the rank of $\sigma_{l+1}^{-1}(Z^{l+1})$, $s_{i}(A)$ denotes the $i$-th singular value of a matrix $A$ and $\sigma_{2},\dots,\sigma_{L}$ is a set of entrywise omeomorphisms with $\sigma_{l}(x)\neq 0$ for all $x\neq 0$.
Constraint (I): representability.
For any realizable sequence of representations $Z^{1},\dots,Z^{L}$, each representation $Z^{l+1}$ must satisfy $Z^{l+1}=\sigma_{l}(W_{l}Z^{l})$ for some weight matrix $W_{l}$. When we reparameterize the representation cost functional in equation 9 in terms of the features embeddings (see equation 12), the constraint (I), by coupling the embeddings $z^{l}$ and $Z^{l+1}$, expresses this representability requirement. We remark that this constraint does not restrict the optimization problem further, as it was shown in (Jacot et al., 2022) that the stationary points of equation 9 are in bijection with the ones of equation 12 with constraint (I).
Constraint (II): effect of nonlinearity.
The second constraint (II) can be directly related to the effect of the nonlinear activations, and it represents exactly the situation for diagonal feature matrices, for which the nonlinearity acts directly on the singular values. The necessity of constraint (II) for the theoretical analysis relies on making it possible to study the alignment of the optimal basis, avoiding any possible singular value escaping to infinity. Interestingly, notice that while in (II) $C$ is an arbitrary positive constant, we observe numerically in Section˜5 that a perfect fit happens for $C=1$, reminiscent of a norm balancing condition on all the layers.
Constraint (III): excluding lower-than rank $K$ solutions.
The third set (III) is introduced here as it arises naturally in the proof of the theorem as a consequence of (II), and it allows us to rule out the presence of global minima of lower rank. This is because, as we will show in the proof in Appendix˜B, all global minima satisfy the equality constraint. We notice that, as a byproduct of Theorem˜4.2, we know that if we get rid of constraint (III), then all global minimizers either satisfy $\operatorname{\mathrm{DNC1}}$ or some layer embedding has rank strictly smaller than $K$. We expect these minimizers to be exceptional in some sense, as they were proven to be optimal in a very particular setting in (Súkeník et al., 2024a). We also remark that the singular values structure is also predicted recursively by Theorem˜4.2, which we will numerically show also in Section˜5.
We highlight here that constraint (II) corresponds to a constraint just on the nonzero singular values of the matrices $Z^{l}$ and $\sigma_{l+1}^{-1}(Z^{l+1})$. Therefore, (II) allows the case in which all feature matrices $Z^{l}$ for $l=1,\dots,L-1$ are full-rank. Moreover, $\mathcal{S}$ is bounded but not compact because of constraint (III), making it nontrivial to prove the existence of a minimizer. The following main result on the minimizers of a constrained version of equation 12 holds:
@@
Let $\sigma_{2},\dots,\sigma_{L}$ be a set of entrywise homeomorphisms and $\sigma_{1}$ an analytic non-polynomial function, with $\sigma_{l}(x)=0$ if and only if $x=0$ for every $l=1,\ldots,L$. Assume $K=n_{L}\leq\dots\leq n_{1}=N\leq 2^{n_{1}-1}$. Let $\mathcal{S}\subset\mathbb{R}^{n_{1}\times N}\times\dots\times\mathbb{R}^{n_{L}\times N}$ be as in Definition˜4.1. Then, for almost every $Z^{1}$, the global minima of the optimization problem
@@
$\bar{Z}^{l}\propto O_{l}\sigma_{l+1}^{-1}(\bar{Z}^{l+1}),\quad O_{l}\in\mathbb{R}^{n_{l}\times n_{l+1}},\;O_{l}^{\top}O_{l}=I,\quad\forall\,l=2,\dots,L-1.$
Proof (sketch).
First, we consider the set of $Z^{1}\in\mathbb{R}^{N\times N}$ such that $\operatorname{rank}(Z^{1})=N$, which is a generic condition as shown in Lemma˜B.3. Now, fixed $Z^{1}$ satisfying this condition we look at the minimizers in the $Z^{2}$ variables, subject to the constraints in Definition˜4.1. Since regularization on the second term $W_{2}$ is not present, there is just a single term involving $Z^{2}$, leading to the following:
$\min_{Z^{2}\in\mathbb{R}^{n_{2}\times N}}\|\sigma_{3}^{-1}(Z^{3})Z^{2,+}\|^{2}$
$\qquad\text{s.t.}\quad\sigma_{3}^{-1}(Z^{3})Z^{2,+}Z^{2}=\sigma_{3}^{-1}(Z^{3})$
$\qquad\qquad\;\sigma_{2}^{-1}(Z^{2})Z^{1,+}Z^{1}=\sigma_{2}^{-1}(Z^{2})$
$\qquad\qquad\;s_{i}(Z^{2})\leq C_{2}s_{i}(\sigma_{3}^{-1}(Z^{3})).$
Notice that, since $Z^{1}$ is invertible, the second representability constraint is trivially satisfied, as $Z^{1,+}Z^{1}=I_{N}$. Therefore, thanks to Lemma˜B.1, we can describe exactly the global minimizers of this optimization problem as a function of $\bar{Z}^{2}\propto O_{2}\sigma_{3}^{-1}(Z^{3})$. By substituting this expression in $\mathscr{L}$, we can do recursively the same for $Z^{3}$, by using the third constraint on the rank of representations. ∎
Remark 4.3.
For almost every $Z^{1}$ is intended in the sense of pushforward Lebesgue measure, i.e., for almost every $(X,W_{1})\in\mathbb{R}^{n_{0}\times N}\times\mathbb{R}^{n_{1}\times N}$.
A rigorous formulation and proof of Theorem˜4.2 involves multiple technical lemmas, and can be found in Appendix˜B.
Theorem˜4.2 can be reinterpreted as follows: if we allow the architecture to collapse already at the first layer, that is, if the kernel of $W_{1}$ can be large enough (i.e., $n_{1}\geq N$), then, under (I)(III), it is optimal to collapse all intermediate layers. Notice that the recursive nature of Theorem˜4.2 describes all the optimal intermediate embeddings $Z^{l}$ recursively as a function of $\sigma_{L}^{-1}(Z^{L})$, both in terms of singular values and singular vectors. As an example, $Z^{L-1}$ has to be proportional to a rotated version of $\sigma_{L}^{-1}(Y)$, and therefore we know singular values and vectors exactly. Therefore, by going backward recursively, we can recover the structure of all of the previous layers.
We emphasize that Theorem˜4.2 can be extended to residual networks under a weaker set of assumptions. Indeed, in that case, we can get rid of all constraints, except the natural one in (I). Given the similarity of the key ideas, we present this result along with its proof in Appendix˜C.
@@
Theorem 4.4 (Benign loss landscape: monotonic paths to $\operatorname{\mathrm{DNC1}}$).
@@
Assume the intermediate widths in $\mathcal{S}$ satisfy the condition $K\leq n_{L}=n_{L-1}=\dots=n_{1}=N$ and that the $\sigma_{l}$’s are analytic. Then, for almost every initial condition $Z(0)\in\mathcal{S}$, there exists a continuous path $\gamma(t)$ such that:
  1. 1.
    $\gamma(0)=Z(0)$ and $\gamma(1)$ is a global minimum of $\mathscr{L}$;
  2. 2.
    $\mathscr{L}\circ\gamma$ is continuous Lebesgue-almost everywhere (except for a finite set of times), with all discontinuities $t_{1},\dots,t_{m}$ removable, i.e., $\underset{t\to t_{i}^{-}}{\lim}\mathscr{L}(\gamma(t))=\underset{t\to t_{i}^{+}}{\lim}\mathscr{L}(\gamma(t))>\mathscr{L}(\gamma(t_{i})),\quad\forall i=1,\dots,m;$
  3. 3.
    for all continuity times $s\geq t$ we have $\mathscr{L}(\gamma(t))\geq\mathscr{L}(\gamma(s)).$
Proof (sketch).
As in Theorem˜4.2, we consider the set of $Z^{1}\in\mathbb{R}^{N\times N}$ such that $\operatorname{rank}(Z^{1})=N$, which is a generic condition as shown in Lemma˜B.3. Now, fixed $Z^{1}$ satisfying this condition we look at the minimizers in the $Z^{2}$ variables. Thanks to Lemma˜B.4, there exists an energy decreasing path from the initial condition to the global minimizer in $Z^{2}$. We recursively concatenate this with the optimal path in $Z^{3}$, where $Z^{2}$ follows the central path dynamics (staying optimal as in (Han et al., 2022)). At the end, we end up with the concatenation of $L$ loss decreasing paths in $\mathcal{S}$, which is therefore globally energy decreasing and connects the initial condition to the constrained global minimizers. ∎